[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed May 10, 2017

Mark DeMaria - NOAA Federal mark.demaria at noaa.gov
Wed May 10 16:23:13 UTC 2017


Monica:

I am not sure. Maybe that was an old ticket. Also, Phil Cragg was here last
week, not Salim.

Mark

On Wed, May 10, 2017 at 12:08 PM, Monica Bozeman - NOAA Federal <
monica.bozeman at noaa.gov> wrote:

> Wait, so what happened with the GOES-E system? I didnt put in a ticket
> recently for that.
>
>
> Monica Bozeman
> Meteorologist/Programmer
> National Hurricane Center
> 301-683-1553 <(301)%20683-1553>
> monica.bozeman at noaa.gov
> <http://www.wdtb.noaa.gov/badges/monica.bozeman>
>
> On Wed, May 10, 2017 at 10:51 AM, TSB Admin <nhc.tsbadmin at noaa.gov> wrote:
>
>> Here is today's summary of NHC computer operations:
>> --- Craig
>>
>> ----------
>>
>> *TAFB:*
>>
>> 1. N-AWIPS issues: Launching the Danger Graphic prompted TAFB forecasters
>> for a password. NCO Sysadmin Brian Maher fixed this [ssh authorized keys]
>> problem within 17 minutes.
>>
>> 2. GFE/NCP issues: The pre-wind speed probabilities did not ingest into
>> GFE/AWIPS for TD One-E/Adrian. AWIPS-II Sysadmin Chris Mello is aware of
>> this problem and is troubleshooting this morning.
>>
>> 3. GFE/NCP issues: The Experimental Offshore Waters Forecast text
>> formatters crashed for the TS Adrian. The problem was attributed to a
>> baseline change to CAVE that requires a ETN [electronic transfer number?]
>> be attached to all tropical hazards in GFE. Shannon White was able to
>> modify our TROPICAL_hazards smart tool to automatically assign an ETN for
>> both basins in GFE and after doing so the text formatters ran fine.
>> *HSU:*
>>
>> 1. ATCF issues: Unable to import satellite fixes (Dvorak classification)
>> from TAFB and SAB since 09/18Z. Senior Hurricane Specialist Mike Brennan
>> determined that this occurred because both TAFB and SAB hadn't renumbered
>> the system from EP90 to EP01 when they submitted their fixes. The storm was
>> renamed TD One-E at dtg (date-time group) = 2017050918 <(201)%20705-0918>
>> and TS Adrian at dtg = 2017051006 <(201)%20705-1006>.
>> *Other items:*
>>
>> 1. Salim Leyva visited NHC last week to restore the signal from the
>> GOES-east antenna. The global imaging system for GOES-east needs to
>> restarted to restore the processing of the local GOES-east satellite data.
>> The sat server currently only processes a few images, then dies. TSB
>> Developer Monica Bozeman has submitted a NCO Help Desk ticket on this issue.
>> *NHC RFCs:*
>>
>>
>>
>
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170510/4926df54/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list